A solution that contains 4.20 g of X in 200 g of chloroform boils at 62.290C. What is the molecular weight of X?


1
Expert's answer
2021-02-01T03:51:23-0500

ΔT = Kbm

m – molal concentration

4.20 / M – moles of X

200 g = 0.2 kg – mass of solvent

So, m = 4.20 / 0.2M


Chloroform boiling point – 61.20С

ΔT = 62.29 – 61.2 = 1.090C


Kb (chloroform) = 3.88


So, 1.09 = 3.88 x 4.20 / 0.2M

M = 74.75 (the molecular weight of X)